AziTech is considering the design of a new CPU for its new model of computer systems for
2021. It is considering choosing between two (2) CPU (CPUA and CPUB) implementations
based on their performance. Both CPU are expected to have the same instruction set
architecture. CPUA has a clock cycle time of 60 ns and CPUB has a clock cycle time of 75
ns. The same number of a particular instruction type is expected to be executed on both CPUs
in order to determine which CPU would executes more instructions. CPUA is able to execute
2MB of instructions in 5*106 clock cycles. CPUB executes the same number of instructions in
3*106 clock cycles.
a) Using the MIPS performance metric, which of the two (2) CPU should be selected to
be implemented in the new computer system. Justify your choice.
b) Compute the execution time for both CPUs. Which CPU is faster?
Answer #1

a)

Here

MIPS = Instruction Count / (Execution time * 10^6)

Execution times of CPUA and CPUB are calculated in section b

MIPS of CPUA = 2 MB / (0.3 Seconds * 10^6) = 2*10^6 / (0.3 *10^6) = 6.67 MIPS

MIPS of CPUB = 2 MB / (0.225 Seconds* 10^6) = 2*10^6 / (0.225 *10^6) = 8.89 MIPS

Its clear from MIPS value of CPUB is higher than CPUA

So CPUB will be selected to implement the new computer system

============================================================

b)   

CPUExecutionTime = CPUClockCycles * ClockCycleTime

Excecution time of CPUA = 5*10^6 * 60*10^-9 = 300 * 10^-3 = 300 milliSeconds = 0.3 Seconds

   Excecution time of CPUB = 3*10^6 * 75*10^-9 = 225 * 10^-3 = 225 milliSeconds = 0.225 Seconds

Generally

Execution TimeA < Execution TimeB

Then, PerformanceA > PerformanceB [ Performance6 = 1 / (Execution Time) ]

Here, Execution Time of CPUB < Execution Time of CPUA, 0.225 Sec < 0.30 Sec

CPUB will be runs faster